OpenStudy (anonymous):

PLEASE HELP WITH THIS VECTORS PROBLEM! a) Write a vector equation and parametric equations of the line through P(3 , 5) and Q(-4 , 7). b) Write the scalar equation of the line with parametric equations: x=6-7t , y=-2+3t Please show all work, Thanks (:

OpenStudy (anonymous):

First work out the vector through those two points P and Q Q(-4, 7) - P(3, 5) vector QP = (3 - (-4) , 5 - 7) = (7, -2) Now that we have a vector we simply write the vector equation the general form of the vector equation is (x,y) = Xo + t(v) (x,y) = (-4,7) + t(7, -2) **** This is the vector equation To obtain parametric equation, you simply break down the above equation for x and y x = -4 + 7t y = 7 -2t

OpenStudy (anonymous):

When does part a) end?

OpenStudy (anonymous):

that is part A, Im not sure about part B

OpenStudy (anonymous):

Oh, so the answer is x = -4 + 7t y = 7 -2t?

OpenStudy (anonymous):

@javawarrior ?

OpenStudy (anonymous):

yes those are the parametric equations

OpenStudy (anonymous):

Thanks :) How do we do part b)?

OpenStudy (anonymous):

From my part of the world, the scalar equation is referred to as the general equation. simply take these equations: x=6-7t y=-2+3t and try to eliminate t The general form the equation should be Ax + By = D

OpenStudy (anonymous):

How do we eliminate t?

OpenStudy (anonymous):

just take a multiple of one equation and minus the other to get rid of t

OpenStudy (anonymous):

Can you take me through it?

OpenStudy (anonymous):

@javawarrior ?

OpenStudy (anonymous):

@lgbasallote can you solve those 2 equations to eliminate T ?